OpenStudy (anonymous):

Multiply and simplify to lowest terms. x^2/x^2+15 *x+7/x

OpenStudy (saifoo.khan):

\[\Large \frac{x^2}{x^2+15} \times \frac{x+7}{x}\]Like this?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

The only "x"s cancel out!

OpenStudy (saifoo.khan):